### v1

# CordViP: Correspondence-based Visuomotor Policy for Dexterous Manipulation in Real-World

Abstract
Achieving human-level dexterity in robots is a key objective in the field of robotic manipulation. Recent advancements in 3D-based imitation learning have shown promising results, providing an effective pathway to achieve this goal. However, obtaining high-quality 3D representations presents two key problems: (1) the quality of point clouds captured by a single-view camera is significantly affected by factors such as camera resolution, positioning, and occlusions caused by the dexterous hand; (2) the global point clouds lack crucial contact information and spatial correspondences, which are necessary for fine-grained dexterous manipulation tasks. To eliminate these limitations, we propose CordViP, a novel framework that constructs and learns correspondences by leveraging the robust 6D pose estimation of objects and robot proprioception. Specifically, we first introduce the interaction-aware point clouds, which establish correspondences between the object and the hand. These point clouds are then used for our pre-training policy, where we also incorporate object-centric contact maps and hand-arm coordination information, effectively capturing both spatial and temporal dynamics. Our method demonstrates exceptional dexterous manipulation capabilities with an average success rate of 90% in four real-world tasks, surpassing other baselines by a large margin. Experimental results also highlight the superior generalization and robustness of CordViP to different objects, viewpoints, and scenarios. Code and videos are available on https://aureleopku.github.io/CordViP.
Figure 1: We propose CordViP, a correspondence-based visuomotor policy for dexterous manipulation in the real world. (a) Left: We present the interaction-aware point clouds, which demonstrate robustness to different viewpoints while establishing correspondences between the object and the hand. (b) Right: Our method achieves promising results across four real-world dexterous manipulation tasks, showcasing exceptional generalization capabilities.
I Introduction
Dexterous manipulation is a fundamental capability in human daily life such as assembling small parts and opening boxes. Achieving human-level dexterity in real-world scenarios is crucial for integrating robots into everyday human activities. Recent advancements in imitation learning have demonstrated significant potential in various robotic manipulation tasks. Some existing methods leverage 2D images as input to directly predict actions [7, 67, 51, 60]. While these vision-based imitation learning approaches are capable of handling a wide range of tasks, they typically demand extensive demonstrations [19, 7, 13, 54] and, at the same time, fail to capture intricate spatial relationships and 3D structures essential for dexterous manipulation [66, 23, 46, 11], limiting their ability to perform complex, fine-grained tasks.
Recent research has increasingly focused on 3D imitation learning, utilizing point clouds and voxels as representations, to help robots perceive 3D environments and reason about spatial relationships [66, 13, 53, 14, 61, 48]. For example, 3D diffusion-based policies, which aim to enhance robots’ ability to perform complex manipulation tasks by providing a more accurate and holistic representation of the environment. These methods for obtaining 3D representations in real-world scenarios often depend on a single-view RGB-D camera to generate point clouds, imposing significant demands on both the camera’s quality and its position. However, the multi-fingered nature of dexterous hands often leads to occlusions during object manipulation, which not only loses the spatial information of the object point clouds but also obscures critical correspondence information for precise manipulation. Besides, some works have explored the use of tactile sensors to enhance contact information [17, 63, 18, 59, 2]. For instance, Wu et al. [59] proposed a canonical representation of 3D tactile data, which is concatenated with other visual information. Although tactile sensors hold promise, their high cost and susceptibility to interference from factors such as temperature fluctuations and electromagnetic fields undermine their stability and limit their practical applicability in real-world scenarios. Amidst all the challenges, leveraging efficient and robust 3D imitation learning to understand spatial information in complex, dynamic environments is crucial for dexterous manipulation.
In this paper, we propose a correspondence-based visuomotor policy (CordViP), which focuses on the spatial and temporal consistency between the manipulated object and the dexterous hand, even under significant occlusions. CordViP separately extracts the 3D representations of robotic hand’s joint structure and the manipulated object, and utilizes the contact map between them as well as arm-hand cooperative motion information to pretrain the observation encoder. This approach enhances the dexterous hand’s ability to understand spatial interactions and better facilitate a range of downstream tasks. Specifically, our framework operates in three phases: (1) We leverage the robust 6D pose estimation of objects combined with the robot’s proprioceptive state to construct interaction-aware point clouds, providing ideal 3D observations that facilitates effective learning and inference for the visuomotor policy. (2) We pretrain the encoder network on the play data we collected to predict the contact map between the dexterous hand and the object, and reconstruct the cooperative relationship between the hand and the arm for further understanding of spatial and temporal correspondences. (3) We utilize the pre-trained encoder to extract the semantic features of 3D point clouds and robot state, which encapsulates spatial consistency, contact dynamics and collaborative information, and then are used as conditions for the diffusion policy to predict actions.
To comprehensively evaluate our proposed CordViP, we conduct extensive real-world experiments on four dexterous tasks: PickPlace, FlipCup, Assembly, and Artimanip. Comparative results demonstrate that our method not only exhibits superior effectiveness, but also achieves remarkable performance with a minor number of expert demonstrations, highlighting its capability to learn efficiently from limited data. Furthermore, we find that CordViP generalizes well to various environmental perturbations, including varying lighting conditions, unseen objects, diverse scenarios, and different camera viewpoints, outperforming other baselines by a large margin.
In summary, our contributions are as follows:
  1. 1.
    We develop a pipeline based on robust 6D pose estimation of objects and robot proprioception, which enables the real-time acquisition of complete 3D representations and semantic flow in real-world environments and can effectively address the challenges posed by occlusions during dexterous manipulation.
  2. 2.
    We propose CordViP, a correspondence-based visuomotor policy that utilizes the contact map and hand-arm coordination information to facilitate the understanding of spatial and temporal consistency.
  3. 3.
    We demonstrate the effectiveness and generalization of our method through a range of real-world experiments using a dexterous hand.
II Related Work
II-A Dexterous Manipulation
Dexterous Manipulation is a long-standing research topic in robotics that aims to give robots the ability to perform delicate operations like humans [5, 1, 42, 62]. Traditional methods often rely on trajectory optimization based on dynamic models to solve operational problems [26, 32, 55], but these methods have limitations in complex tasks because they simplify the contact dynamics and are difficult to deal with uncertainties in dynamic environments. In contrast, Reinforcement Learning (RL) does not rely on accurate physical models but learns operational policies through interaction with the environment, which is highly adaptable. RL has achieved remarkable results in many dexterous manipulation tasks, such as object reorientation [3, 40, 36, 20] and sequential manipulation [6, 15]. However, RL methods often suffer from several challenges, such as the need for extensive reward engineering and system design, as well as limited generalization to unseen scenarios. Additionally, while Sim-to-Real is a common technique employed in RL, the gap between simulations and the real world degrades the performance of the policies once the models are transferred into real robots [68]. Imitation learning (IL), as another effective learning method, can quickly learn effective control policies by imitating expert demonstrations [64, 51]. In this work, we propose a correspondence-based visual imitation learning policy that utilizes spatial information between various components, enabling the acquisition of complex skills with a minimal number of expert demonstrations.
II-B Imitation Learning
Imitation learning (IL) allows a robot to directly learn from experts. Behavioral Cloning (BC) is one of the simplest imitation learning algorithms, which treats the problem of learning behavior as a supervised learning task [37]. The modeling methods commonly used in traditional BC, such as MSE, discretization [29], and K-Means [16], have limitations when modeling complex action distributions. They fail to effectively capture the diversity and nuances of human behavior [35]. Over the past few years, diffusion models have emerged as a new modeling approach in BC, becoming powerful tools that enable robots to learn from demonstrations, handle uncertainty, and perform complex multi-step tasks with precision. From the early applications of DDPMs to the recent innovations in BESO [44], OCTO [49], and CrossFormers [10], these models have continually pushed the boundaries of what’s possible in robotic behavior generation. While traditional BC policies typically rely on 2D image-based representations [7, 67, 51, 60, 28], recent advancements have extended imitation learning to 3D visual representations [66, 4, 13, 52, 53, 30]. These 3D approaches provide a more comprehensive understanding of spatial relationships and 3D structures, further enhancing robotic behavior learning.
II-C Correspondence Learning
Correspondence refers to the relationship or alignment between different entities or components, with the aim of establishing meaningful connections. Correspondence learning has been shown to improve performance in various robotic tasks, including grasping [33, 9], perception [27, 4], pose estimation [21] and garment manipulation [58]. In this paper, correspondence specifically refers to the alignment between hand-object spatial interaction and hand-arm temporal coordination. By incorporating correspondence, we enhance feature extraction capabilities, thereby enabling more accurate and coordinated movements in downstream tasks.
III Method
The overview of our framework is shown in Figure 2, which operates in three phases: (1) Interaction-aware generation of 3D point clouds. We acquire relatively accurate and complete 3D observations during real-world dexterous manipulation tasks even under significant occlusions, as described in III-B. (2) Contact and coordination-enhanced feature extraction. By leveraging large-scale play data and incorporating contact maps and hand-arm coordination, we improve spatial interaction perception and capture cooperative motion features, detailed in III-C. (3) Correspondence-based diffusion policy. The pre-trained encoder is used to extract 3D representations, which guide the training of a visuomotor policy, as outlined in III-D.
Figure 2: Overview Framework (a) We first employ TripoSR to generate the initial object point cloud and FoundationPose to estimate the 6D pose of the object. In parallel, the hand point cloud is generated based on the robot’s state. They are combined to construct interaction-aware point clouds, which demonstrate robustness to viewpoint variations. (b) During the pre-training phase, the generated point cloud data, combined with the robot’s proprioceptive information, is utilized to enhance spatial understanding and interaction modeling. (c) The pre-trained encoder is subsequently integrated into an imitation learning framework to facilitate downstream tasks in dexterous manipulation.
III-A Problem Formulation
We formulate our problem as learning a visuomotor policy $\pi:\mathcal{O}\mapsto\mathcal{A}$ from expert demonstrations of the form of $\{(o_{1},a_{1}),(o_{2},a_{2}),\dots,(o_{n},a_{n})\}$, where $\mathcal{O}$ represents the robot’s observations and $\mathcal{A}$ represents the corresponding actions, allowing the robot to generalize beyond the training data distribution. In our approach, each observation $o_{t}$ is composed of the object’s point cloud $\mathbf{P}_{obj}$, the hand’s point cloud $\mathbf{P}_{hand}$, and the robot’s joint states, including a 6-Dof arm and 16-Dof hand configuration. Unlike previous works that rely on global point clouds for 3D feature extraction, our approach prioritizes capturing the individual information of the arm, hand, and object throughout the manipulation process. As a result, CordViP not only effectively addresses occlusion challenges during dexterous manipulation but also significantly improves the model’s ability to comprehend spatial interactions and collaborative dynamics. Furthermore, leveraging these observations, we compute contact maps between the robotic hand and the manipulated objects, as well as capture collaborative interaction information between the arm and hand. These elements are critical for modeling spatial and temporal relationships.
III-B Interaction-aware Generation of 3D Point Clouds
Motivated by the superior generalization and efficiency of the 3D-based diffusion policy[66, 53, 65, 4], the key intuition behind our solution is to focus on the interactions between the hand and the manipulated object in 3D space. Although intuitively reasonable, achieving this goal is challenging in practice. On the one hand, real-world point cloud data, typically captured using stereo cameras or low-cost RGB-D scanners, suffers from geometric and semantic loss due to factors such as light reflection, material transparency, and limitations of sensor resolution and viewing angle. On the other hand, during dexterous manipulation with multi-fingered hands, occlusions frequently occur, resulting in the loss of critical contact and interaction information, which is vital for precise and effective manipulation. To this end, we propose the interaction-aware generation of 3D point clouds, enabling the reconstruction of crucial spatial information.
Real-to-Sim for Digital Twin Generation. To achieve the goal of obtaining a complete and accurate static point cloud of the manipulated object, we aim to reconstruct the digital twin from a single-view image. Referring to the approach TripoSR [50], we implement a 3D generation technique, which utilizes its strong priors and broad understanding of visual concepts in the 3D world to generate 3D digital assets. To ensure the accuracy of point cloud flow tracking, we maintain consistency between the geometric and material properties of the reconstructed assets and their real-world counterparts. Subsequently, we uniformly sample points on the surface of the generated digital twin to obtain the initial 3D point clouds, providing a robust and accurate initial observation for both 3D spatial perception and pose tracking.
Pose-Driven Point Cloud Tracking. We have successfully obtained the initial point cloud of the object. However, tracking the object’s point cloud in complex, real-world environments poses a significant challenge, particularly in scenarios with severe occlusions. To overcome this, we leverage foundation models to ensure precise pose estimation of the manipulated object, thereby enhancing point cloud tracking. Specifically, we first employ the Segment Anything model [25] to extract the mask of the manipulated object. This is combined with a digital twin generated via real-to-sim techniques, enabling us to use FoundationPose [57] for accurate pose estimation. Using the center of the acquired point cloud as the origin, we then apply the estimated pose to transform the entire point cloud, aligning it with the desired coordinate frame.
Robot Point Cloud Forward Kinematics Model. To obtain the point cloud of the dexterous hand, we develop the robot point cloud forward kinematics model $\mathcal{F}_{pc}$. We first parse the URDF file to identify individual links of the robot system, and uniformly sample point clouds on the surface for each link. The robot system model is constructed with point clouds denoted as $\{P_{\ell_{i}}\}_{i=1}^{N_{\ell}}$, where $N_{\ell}$ is the number of links. Following Wei et al. [56], we designed the robot point cloud forward kinematics model that maps any joint configuration to the corresponding pose of the point cloud. In order to focus on the interaction between the hand and the object, we ignore the point cloud of the robotic arm. Therefore, the 3D observation of the hand $\mathbf{P}_{hand}\in\mathbb{R}^{N_{\mathcal{P}}\times 3}$ is defined as:
$\mathbf{P}_{hand}=\mathcal{F}_{pc}(q,\{P_{\ell_{i}}\}_{i=1}^{N_{\ell}}),$ (1)
where $N_{\mathcal{P}}$ represents the downsampled size of the point cloud, which is set to 1024 in practice. Through our practical evaluations, the point cloud forward kinematics model operates at 8Hz, a frequency that is sufficiently high for our experimental setup, where inference is conducted at 5Hz.
As shown in Figure 3, interaction-aware point clouds significantly enhance the quality of 3D observations compared to RGB-D synthesized point clouds, while demonstrating strong robustness to different viewpoints.
Figure 3: Point Clouds Comparison. We present point clouds of two methods under three different viewpoints. Notably, for better visualization, we have applied color information to the point clouds. However, color information is not used in the policy learning.
III-C Contact and Coordination-Enhanced Feature Extraction
Interaction-aware Generation of 3D Point Clouds provides us with accurate and complete point cloud observation. However, fine-grained dexterous manipulation often requires more detailed information. pre-training based on contact and coordination, in the context, can encourage the encoder to learn the intrinsic structures, facilitating more effective feature extraction, and promoting the learning of various downstream tasks.
Contact Map Synthesis. A contact map serves as a critical piece of information in dexterous manipulation tasks, which captures the interaction between the hand and the object. Given the complete point clouds, we first calculate object-centric contact map $\mathcal{C}$ as the normalized distances from the object’s surface point to the hand surface. Given the point clouds of the object $\mathbf{P}_{obj}$ and the point cloud of the hand $\mathbf{P}_{hand}$, the aligned distance $\mathcal{D_{(O,H)}}$ between each point $v_{o}$ on the object surface and the surface of the dexterous hand is defined as follows:
$\mathcal{D_{(O,H)}}=\min_{v_{h}\in\mathbf{P}_{hand}}e^{\gamma(1-|<v_{h}-v_{o},n_{o}>|)}\|v_{o}-v_{h}\|_{2},$ (2)
where $n_{o}$ is the surface normal of the object, which is computed using the K-Nearest Neighbors methods [8] by considering the local geometric properties of the point cloud. $\gamma$ is the scaling factor, which is empirically set to 1.
Based on the aligned distance, we compute the contact map following Jiang et al. [24]:
$\mathcal{C}=1-2(\text{Sigmod}(\theta\cdot\mathcal{{D_{(O,H)}}})-0.5)$ (3)
where $\theta$ is the scaling factor and each point’s contact value $c_{i}\in\mathcal{C}$ is bounded within the [0, 1] range.
Point Cloud Feature Extraction. We employ two encoders with identical architecture to extract point cloud embeddings, denoted as $f_{\theta_{O}}(\mathbf{P}_{obj})$ and $f_{\theta_{H}}(\mathbf{P}_{hand})$, for the object and hand, respectively. In detail, we adopt PointNet [38] as the point cloud encoder, which excels in capturing local structures and integrating global information. To establish correspondences between the hand and the object features, we apply two multi-head cross-attention transformers $g_{\theta_{O}}$, $g_{\theta_{H}}$ to fuse their respective embeddings, which maps the hand and object features to two sets of aligned representations, denoted as $\phi^{R}$ and $\phi^{O}$:
$\begin{split}\phi^{H}&=g_{\theta_{H}}(f_{\theta_{H}}(\mathbf{P}_{hand}),f_{\theta_{O}}(\mathbf{P}_{obj}))+f_{\theta_{H}}(\mathbf{P}_{hand})\\ \phi^{O}&=g_{\theta_{O}}(f_{\theta_{O}}(\mathbf{P}_{obj}),f_{\theta_{H}}(\mathbf{P}_{hand}))+f_{\theta_{O}}(\mathbf{P}_{obj})\end{split}$ (4)
To help the point cloud encoder learn the intrinsic features, we designed a contact map prediction task. Since 3D point cloud observations implicitly contain contact information, we utilize a three-layer MLP to predict the object-centric contact map $\mathcal{C}^{pred}$ given the point cloud observations of both the hand and the object. The MSE loss is calculated between $\mathcal{C}$ and $\mathcal{C}^{pred}$. This pre-training approach enables the encoder to learn the interactions and relationships within the environment.
Hand-Arm Coordination Enhancement. To help the robot system learn the features of hand-arm coordination, we also propose a correspondence-based design for action prediction. The arm and hand states are first projected into vectors of identical dimensionality through a linear layer, after which the same cross-attention transformers are employed to establish correspondences between the hand and the arm. We predict the action sequence of the robot arm based on the point clouds and the state of the hand. Similarly, we also predict the action sequence of the hand using point clouds and the arm state. We use MSE loss to compute the loss between the reconstructed and original action. By further predicting the action sequence respectively, the encoder is able to learn intrinsic features of motion and capture collaborative dynamics.
Given the aforementioned losses, our overall training objective during the pre-training phase is defined as:
$\underset{\mathcal{E}}{\min}\mathcal{L}=\mathcal{L}_{contact}+\lambda\mathcal{L}_{coordination},$ (5)
where $\mathcal{E}$ represents the encoder of the observation, and $\lambda$ is a hyperparameter that controls the relative strengths of the losses.
III-D Correspondence-based Diffusion Policy
After obtaining the pre-trained encoder, we utilize an imitation learning framework to learn visuomotor policy for dexterous manipulation tasks. Specifically, we adopt conditional denoising diffusion model [22, 7, 35] as our backbone, which conditions on 3D visual features $\phi^{H,O}$ and robot states features $\psi^{A,H}$. Beginning with a Gaussian noise $A^{K}$, the denoising network $\epsilon_{\theta}$ performs $k$ iterations to gradually denoise $A^{K}$ into the noise-free action $A^{0}$:
$A^{k-1}=\alpha_{k}(A^{k}-\gamma_{k}\varepsilon_{\theta}(\phi^{H,O},\psi^{A,H},A^{k},k))+\sigma_{k}\mathcal{N}(0,I),$ (6)
where $\mathcal{N}(0,I)$ is Gaussian noise, $\alpha_{k}$, $\gamma_{k}$ and $\sigma_{k}$ are functions of $k$, determined by the noise scheduler. This formulation allows the model to capture the distribution of action without the cost of inferring future states.
We use the DDIM scheduler [47] to accelerate the inference speed in real-world experiments. The training objective is to predict the noise added to the original data:
$\mathcal{L}=MSE(\varepsilon^{k},\varepsilon_{\theta}(\bar{\alpha_{k}}A^{0}+\bar{\beta}\varepsilon^{k},\phi^{H,O},\psi^{A,H},k))$ (7)
Unlike the original diffusion-based policy, we incorporate consistency-related features as a condition for policy learning using the pre-trained encoder and fine-tuning the encoder during downstream tasks training. This encoder implicitly extracts contact and coordination information, thereby enhancing the policy’s understanding of spatial relationships.
IV Experiments
We conduct comprehensive real-world experiments to answer the following questions:
  • To what extent can our framework promote the learning of the visuomotor policy for dexterous manipulation across diverse real-world scenarios (Section IV-B)?
  • How promising is CordViP in terms of sample efficiency and generalization capability (Section IV-C, IV-D)?
  • What role does each of the system components play in enhancing its overall performance (Section IV-E, IV-F)?
IV-A Experiment Setup
Robot System Setup. As shown in Figure 4, our system consists of a 6-Dof UR5 robot arm and a 16-Dof Leap Hand [45] with four fingers. A single Intel Realsense L515 RGBD camera is mounted on the side of the robot to capture visual observation.
Figure 4: Real robot system. Our system consists of a Leap Hand and a UR5 Arm, with a fixed Realsense L515 camera employed to capture visual observation. The Realsense D435 camera is only used for data collection during teleoperation and is not involved in the policy learning.
Tasks. We evaluate our approach on four dexterous fine-grained manipulation tasks, as shown in Figure 5. The episode length of each task will be limited to a maximum of 380 steps and each task is evaluated with 20 trials. We now briefly describe our tasks:
  1. 1.
    PickPlace. The Leaphand picks up a toy chicken and places it into a blue bowl.
  2. 2.
    FlipCup. The Leaphand reaches a cup lying on the table, lifts it up, and then rotates the wrist to position the cup upright on the table.
  3. 3.
    Assembly. The Leaphand reaches and grasps a cylindrical cup, and insert it into a kettle.
  4. 4.
    ArtiManip. The Leaphand lifts the lid of a box using its thumb and gently opens it, which involves the manipulation of the articulated objects.
Figure 5: Visualization of four dexterous manipulation tasks, with the right side showing the end state.
For each task, we incorporate a certain level of randomization to ensure that the policy learns the task-specific features rather than fitting the trajectory. As shown in Figure 6, the objects are randomly placed within the red rectangular region. More details are provided in Appendix -A.
Figure 6: Randomization of Object Positions. The red rectangles mark the range of positions of manipulated objects. For PickPlace and FlipCup, both the toy chicken and the cup are randomly rotated within a certain range.
Expert demonstrations. Our training demonstrations are collected by human teleoperation. Since our tasks place significant emphasis on coordination and coherence between the hand and the arm, we employ a vision-based approach to teleoperate the robot. Specifically, we use HaMeR [34] to track human hand pose with a single Realsense D435 camera and use Anyteleop [43] framework to retarget the robot system. The robotic arm is controlled through the 6-Dof end-effector’s pose while the robotic hand is controlled through the retargeted hand joint positions. We collect 50 demonstrations for each task, with all expert demonstrations recorded at a frequency of 5Hz.
Baselines. We compare our method with six state-of-the-art imitation learning algorithms, i.e., three vision-based policies(BCRNN [31], ACT [67], DP [7]) and three 3D-based policies(BCRNN+3D, ACT+3D, DP3 [66]). BCRNN+3D and ACT+3D are variants of BCRNN and ACT, respectively, in which the image input is substituted with point clouds, encoded using PointNet. In fact, we also explore a more novel imitation learning framework, RISE [53]. However, we believe this method is not suitable for dexterous hand manipulation tasks, as it focuses solely on 3D observation and lacks crucial joint state information. Further details on the baseline choices and result analysis can be found in the Appendix -C. We emphasize that the resolution of both the image and depth for all 2D and 3D baseline methods is kept consistent, and the point clouds are synthesized from the RGBD camera.
IV-B Effectiveness
The results of the effectiveness experiments are given in Table I. Our proposed CordViP maintains a completion rate of over 85% across four tasks, significantly outperforming the other baselines. BCRNN and its 3D variant perform poorly in all real-world tasks. While the dexterous hand can locate and reach the object, a substantial finger jitter is observed during the grasping and contact-rich manipulation phases. The image-based policies, ACT and DP, achieve good performance in the flipcup and pickplace tasks. However, they struggle with the assembly task. This is mainly due to the significant occlusions during manipulation, where these image-based policies fail to effectively leverage spatial information. Compared to ACT, ACT+3D shows superior results in all tasks, highlighting the crucial importance of geometric structure and spatial information in policy learning.
Interestingly, we discover that the recently proposed DP3 underperforms relative to DP in most of our tasks, which aligns with the findings of Wang et al. [53]. Although several studies have demonstrated the effectiveness of DP3 in simulation and relatively ideal real-world scenarios, such as those with multiple viewpoints, we observe that the quality of the point clouds has a significant impact on DP3’s performance. Further details will be discussed in Section IV-E. In contrast, our method exhibits remarkable robustness to the quality of point clouds. By establishing correspondences between the dexterous hand and objects, CordViP facilitates more effective perception of spatial and interaction information.
Inference Speed. CordViP also attains an efficient inference speed, as it eliminates the need for the farthest point sampling of point clouds while utilizing compact 3D representations. We evaluate the inference speed of CordViP compared with DP3, CordViP reaches a maximum of 12.84 FPS, surpassing DP3’s 11.79 FPS. This highlights that our approach not only achieves enhanced performance but also maintains low computational overhead during inference.
TABLE I: Main results of four real-world tasks. Each experiment is evaluated with 20 trials.
Method PickPlace FlipCup Assembly ArtiManip Avg
BCRNN 0% 0% 0% 5% 1%
BCRNN+3D 0% 5% 0% 10% 4%
ACT 45% 70% 25% 65% 51%
ACT+3D 70% 80% 35% 70% 64%
DP 55% 65% 20% 35% 44%
DP3 30% 20% 0% 40% 23%
CordViP(Ours) 85% 90% 90% 95% 90%
IV-C Efficiency
The number of expert demonstrations plays a crucial role in the performance of imitation learning. To access the learning efficiency, we train ACT, DP, DP3 and our proposed CordViP on four tasks using varying quantities of expert demonstrations. As illustrated in Figure 7, CordViP exhibits superior performance, achieving higher accuracy with fewer demonstrations. Remarkably, even with just 10 demonstrations, CordViP effectively establishes correspondences, extracts spatial and geometric features, and maintains a high success rate. Furthermore, when provided with a sufficient number of demonstrations, CordViP also demonstrates excellent manipulation capabilities.
Figure 7: Experimental results of efficiency. We train ACT, DP, DP3, and CordViP on the PickPlace and FlipCup tasks with an increasing number of demonstrations.
IV-D Generalization
Besides the remarkable effectiveness and efficiency, CordViP also showcases excellent generalization capabilities in real-world dexterous manipulation tasks. In this section, we comprehensively investigate its generalization abilities across four aspects, as detailed below.
Generalize to different lighting conditions. We set up three different lighting conditions. As shown in Table II, our proposed CordViP, based on 3D point clouds without color information, demonstrates a strong generalizability against various lighting conditions. In contrast, diffusion policy fails to complete tasks under most of these conditions. These image-based policies often rely on data augmentation to enhance generalization, which may lead to instability during training.
TABLE II: Generalization results on different lighting conditions. We evaluate the policy under three lighting scenarios: dim light(dim), white light(white) and colored lighting(colored).
Method PickPlace FlipCup
Dim White Colored Dim White Colored
DP 0% 25% 0% 0% 50% 0%
ACT+3D 65% 55% 70% 75% 80% 75%
CordViP(Ours) 80% 85% 80% 80% 85% 90%
Generalize to different scenarios. CordViP relies on a robust 6D pose estimator to track objects and build interaction-aware 3D point clouds, which enables our method to generalize to different scenarios. The results presented in Table III show that DP is highly sensitive to visual variations, while ACT+3D performs well in scenarios with different visual appearances. However, 3D-based policy struggles in more challenging, cluttered scenes, as randomly placed objects can interfere with the synthesis of point clouds. In contrast, our method effectively focuses on the manipulation of subjects and objects, establishing spatial correspondences, and has demonstrated strong generalization to different scenarios.
TABLE III: Generalization results in diverse scenarios, including varying visual appearances and challenging cluttered environments.
Method PickPlace FlipCup
Scene-1 Scene-2 Cluttered Scene-1 Scene-2 Cluttered
DP 0% 0% 0% 0% 0% 5%
ACT+3D 50% 70% 5% 80% 70% 20%
CordViP(Ours) 70% 80% 75% 75% 80% 85%
Generalize to unseen objects. To validate policies’ generalization to different objects, we test them on the PickPlace and FlipCup tasks using three unseen objects, each varying in color, shape, and dynamics. As shown in Table IV, DP shows poor generalization to unseen objects, while ACT+3D shows a certain level of generalization ability. However, ACT+3D struggles to adjust effectively and promptly when confronted with objects that exhibit significant differences. CordViP demonstrates strong generalization ability when handling unseen objects, making fine-grained adjustments by establishing spatial and temporal correspondences.
TABLE IV: Generalization results on unseen objects. For PickPlace and FlipCup tasks, we chose three previously unseen objects, varying in color, shape, and dynamics.
Method PickPlace FlipCup
Object-1 Object-2 Object-3 Object-1 Object-2 Object-3
DP 0% 5% 0% 15% 35% 0%
ACT+3D 55% 85% 75% 55% 40% 75%
CordViP(Ours) 75% 85% 90% 60% 45% 75%
Generalize to different viewpoints. Achieving generalization to different viewpoints presents a more significant challenge, as the expert demonstrations are all captured using a fixed camera. We evaluate DP, ACT+3D, and our method with three different camera viewpoints. For 3D-based methods, since the point clouds are represented in the world coordinate system, the camera is recalibrated for each viewpoint. As shown in Table V, the image-based diffusion policy is highly sensitive to camera viewpoints and completely fails across all three camera views. ACT+3D, which leverages 3D information, demonstrates a certain level of generalization to minor viewpoint changes. However, due to the limitations of the camera viewpoint, the synthesized point clouds can only capture partial spatial information, making it hard to handle significant changes in camera views. Our method, which generates interaction-aware point clouds, is capable of handling significant variations in camera perspectives.
TABLE V: Generalization results to different viewpoints.
Method PickPlace FlipCup
View-1 View-2 View-3 View-1 View-2 View-3
DP 0% 0% 0% 0% 0% 0%
ACT+3D 20% 45% 60% 0% 40% 30%
CordViP(Ours) 80% 75% 65% 60% 70% 80%
TABLE VI: Ablation experiments on the Effectiveness of different components. DP3 + Interaction-aware PC refers to using the Interaction-aware point clouds as the visual input for DP3. W/o. contact and coordination pretrain means that the encoder was not pre-trained with contact and coordination data.
Ablation PickPlace FlipCup Assembly ArtiManip Avg
DP3 + Interaction-aware PC 65% 75% 60% 85% 71%
CordViP w/o. contact and coordination pretrain 75% 80% 75% 85% 79%
CordViP w/o. contact pretrain 85% 75% 85% 90% 84%
CordViP w/o. coordination pretrain 80% 85% 80% 85% 83%
CordViP(Ours) 85% 90% 90% 95% 90%
((a)) Case1: FlipCup.
((b)) Case2: Assembly.
Figure 8: Failure case. (a) Case 1 is a failure case from the Flip Cup task; (b) Case2 is an example from the Assembly task, where after a failure to insert the cup, the policy made adjustments, corrected the orientation of the cup, and successfully completed the task.
IV-E Ablations
Effectiveness of components. We conduct a series of ablation experiments on four real-world tasks to evaluate the effectiveness of different components in our method. As shown in Table VI, our results reveal the critical importance of contact and coordination information for learning dexterous manipulation policy. Specifically, for contact-rich tasks such as Flip-Cup, the absence of pre-training on contact information significantly reduces success rates. This highlights the essential role of contact in understanding physical interactions between the robotic hand and objects. On the other hand, for tasks that require continuous motion of both the arm and hand, such as Assembly and ArtiManip, incorporating coordination information notably enhances the policy’s ability to execute complex movement patterns. These findings emphasize the necessity of both contact and coordination information to achieve robust manipulation.
We further observe that merging the Interaction-aware point clouds into a unified representation and providing it as the 3D input to DP3 leads to a noticeable improvement in the success rates across all tasks. This also suggests that, compared to ACT+3D, DP3 places higher demands on the quality of the point clouds, and is less robust to noise and occlusion in real-world scenarios. In our settings, we do not carefully position the camera to minimize occlusion of the dexterous hand, which demonstrates the robustness of our method to different viewpoints.
Choice of point cloud encoders. We use PointNet [38], PointNet++ [39], and PointNeXt [41] as point cloud encoders. As shown in Table VII, PointNet outperforms the other encoders in success rates. We also freeze the point cloud encoder during the training phase of the correspondence-based diffusion policy. The results indicate that fine-tuning further enhances the learning of downstream tasks.
TABLE VII: Ablation experiments on point cloud encoders. * indicates that the encoder is frozen during the training phase of the correspondence-based diffusion policy.
Encoders PickPlace FlipCup Assembly ArtiManip Avg
PointNet 85% 90% 90% 95% 90%
PointNet++ 5% 30% 30% 40% 26%
PointNeXt 0% 0% 0% 0% 0%
PointNet* 70% 90% 75% 80% 79%
PointNet++* 5% 15% 5% 30% 14%
IV-F Transferability
Our proposed CordViP can be easily transferred to different backbones. Following Zhao et al. [67], we choose a Transformer-based CVAE as the decision backbone for the policy, which outputs a sequence of actions. We replace the 2D image input with our Interaction-aware point clouds and use the pre-trained encoder to establish correspondences between various components. Table VIII illustrates that CordViP achieves excellent performance on both diffusion-based and transformer-based backbones, showcasing strong transferability.
TABLE VIII: Transferability to different backbone.
Method PickPlace FlipCup Assembly ArtiManip
Diffusion-based 85% 90% 90% 95%
Transformer-based 90% 75% 90% 80%
IV-G Failure Case Analysis
In this section, we analyze some failure cases of CordViP. We take the flipcup task as an example and visualize the results in Figure 8(a). We observe that while CordViP generally achieves good localization and grasping, the hand occasionally hovers in place after wrist rotation. One possible reason is that, during the collection of expert demonstrations, we deliberately reduced wrist speed, resulting in slight jitters that led the model to learn suboptimal solutions. A potential solution could be to expand the policy’s values of $horizon$ and $n\_action\_steps$.
Nevertheless, we also observe that our model possesses the ability to automatically correct some failure scenarios and recover from them. In the assembly task, inserting the cylindrical cup into the kettle is a delicate operation that relies on strong 3D spatial reasoning capabilities. As shown in Figure 8(b), CordViP is able to make timely adjustments, such as when the cup is stuck at the opening of the kettle. It can quickly sense the state of the object and adjust the joint pose to successfully complete the assembly task.
V Conclusions and Limitations
In this paper, we present CordViP, a novel framework that learns correspondence-based visuomotor policy for dexterous manipulation in the real world. First, we leverage robust 6D pose estimation of objects and robot proprioception to obtain interaction-aware point clouds, enhancing the quality of the point clouds and addressing the issues caused by occlusions from the dexterous hand. Second, we use object-centric contact map and coordination information to design a pre-training task that effectively establishes spatial and temporal correspondences. Finally, the features obtained from the pre-trained encoder are used as conditions to train a visuomotor policy. CordViP significantly outperforms state-of-the-art 2D and 3D baselines on four real-world dexterous manipulation tasks, demonstrating highly competitive performance in both effectiveness and efficiency.
Limitations. Despite the exceptional performance demonstrated by CordViP, there are still certain limitations that could be explored in future work. First, our method struggles to accurately estimate the 6D pose of deformable objects due to the limited expressive capacity of FoundationPose. Second, the accuracy of digital twin modeling can significantly impact pose tracking and the quality of initial object point clouds, which are critical for fine-grained dexterous manipulation. One potential solution could be the inclusion of an additional wrist-mounted camera. We leave further exploration of these possibilities for future work.